OpenStudy (amy0799):

Consider the function below f(x)= 8x^4-28x+9 Find the critical numbers of f (if any). Find the open interval on which the function is decreasing. Find the open interval on which the function is increasing Apply the First Derivative Test to identify the relative extremum.

OpenStudy (lochana):

you should factorize your equation first of all. and then draw the graph. you will able say increasing and decreasing intervals.

OpenStudy (amy0799):

i didnt know how to solve for f'(x)=4(8x^3-7)=0

OpenStudy (welshfella):

are you familiar with calculus - derivatives?

OpenStudy (amy0799):

yea im taking it

OpenStudy (welshfella):

ok so can you find the derivative of this function?

OpenStudy (lochana):

give me few minutes, i will send you my solution

OpenStudy (amy0799):

f'(x)=32x^3-28

OpenStudy (welshfella):

ok to find critical values equate this to zero and solve for x

OpenStudy (amy0799):

4(8x^3-7)=0 This is where i got stuck. How do i solve the the cube root?

OpenStudy (welshfella):

x^3 = 28/32 use your calculator

OpenStudy (welshfella):

- its the same as cube root of 7/8

OpenStudy (amy0799):

so it's 0.956 and -0.956?

OpenStudy (welshfella):

not not negative because x^3 is positive

OpenStudy (amy0799):

oh ok. so there's only 1 critical point?

OpenStudy (welshfella):

yes If you have a graphic calculator you''l see that this is a minimum The question asks you to identify the realive extrema using first derivative test

OpenStudy (welshfella):

You need to find the slope of the curve each side of the minimum point by finding the sign of f' at values just less than 0.956 and just above I suggest you use 0.95 and 0.96

OpenStudy (welshfella):

if we go from negative to postiuve then its a minimum - positive to negative is a maximum.

OpenStudy (welshfella):



OpenStudy (welshfella):

the critical value of f is the value of f(x) when x = 0.956

OpenStudy (amy0799):

i still dont understand how u find the relative extremum

OpenStudy (welshfella):

ACtually we've done things the wrong way around here. The first part is to draw the graph of f to answer the first 3 parts

OpenStudy (welshfella):

Do you understand why f' = 0 at the extrema?

OpenStudy (amy0799):

is it because the slope is 0?

OpenStudy (welshfella):

Yes f' gives the value of the slope at any point on the graph and at the extrema the slope is horizontal

OpenStudy (amy0799):

what's the next step?

OpenStudy (welshfella):

to identify the nature of the extremium you do the first derivative test f'(x) = 32x^3 - 28 at x = 0.95 f' = 32(0.95)^3 - 28 = -0.564 - showing a negative slope just before the extremium at x = 0.96 f' = 32(0.96)^3 - 28 try working this one out If its positive it shows that the point is a minimum

OpenStudy (amy0799):

0.311 a positive slope

OpenStudy (welshfella):

right that proves its a minimum

OpenStudy (amy0799):

so the point is (0.965,0.311)?

OpenStudy (welshfella):

no the point is (0.956, f(0.956)) plug x = 0.956 into the function 8x^4 - 28x + 9 this is the value of f at the extremium

OpenStudy (welshfella):

you should get -11 point something

OpenStudy (amy0799):

ok so the relative extremum is ((0.956, -11.086)?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

and the intervals are decreasing (-INF, -11.086) - i'll let you workout the interval for increasing f

OpenStudy (amy0799):

shouldn't it be (INF, 0.956)?

OpenStudy (welshfella):

Hmmm the values of x you mean Yes I think you are right

OpenStudy (welshfella):

so decreasing is (-INF,0.956)

OpenStudy (welshfella):

yes my mistake

OpenStudy (amy0799):

i thought it would be positive infinity

OpenStudy (welshfella):

No the values are all less than 0.956 so it does back to negative infinty

OpenStudy (welshfella):

the increasing interval will be (0.956, INF)

OpenStudy (welshfella):

Note the round parentheses because at 0.956 the slope is horizontal

OpenStudy (amy0799):

i submitted my work but i got the critical numbers wong :/

OpenStudy (welshfella):

did they give you the numbers?